# In Situ AI Prototyping: Infusing Multimodal Prompts into Mobile Settings with MobileMaker

### v1

@@ IV-C Prototype Testing Experience on Mobile
Prototype creators can share their prototype via URL, which enables team members or other users to try it out in the wild (D2). When running a prototype on a mobile device, each set of user inputs, model output, and user feedback is saved as a test case. Prototype creators can review all collected test cases in a testing dashboard, allowing them to quickly explore use cases and discover unexpected edge cases.
@@ IV-C1 Dynamic Prototype Revision
When users experience unexpected outputs or prototype behavior, or they have feedback on how the prototype could be improved to better meet their needs, they can leverage the NL revision feature (sec. IV-B2) to generate a revised version of the prototype; this can be sent back to the prototype creator as a suggested revision. Each suggested revision includes the user’s revision request, the updated JSON prototype specification, and the latest test case generated with the revised prototype prior to submission. After revising, users can choose to either continue testing with the modified prototype or revert to the original specification, allowing for both continued iteration on a specific concept and exploration of multiple orthogonal ideas.
@@ IV-C2 Prototype Revision Dashboard
Prototype creators can use the testing dashboard to review and examine all proposed revisions, as well as dynamically try out each revision by applying its changes. When a revision is applied, MobileMaker renders the layout and functionality of its JSON prototype specification in the mobile phone preview, then populates the preview with inputs and outputs from the associated test case that the original tester generated and experienced. This enables prototype creators to quickly understand and explore the impact of each revision on their prototype, significantly enhancing and accelerating the prototype creator’s design loop (D3).
@@ V Exploratory User Study
To understand how MobileMaker impacts the feedback users provide, we conducted a within-subjects study where participants gave feedback on two prototypes, one in each study condition. In the MobileMaker condition, users experienced a MobileMaker prototype in the wild, walking and taking photos of their environment to experience the prototype on their phones. In the baseline (desktop) condition, they clicked through Figma mockups demonstrating the basic interaction, then searched for images online to test the LLM prompt that would power the prototype. We chose this setup based on formative studies with designers, who indicated that these are typical prototyping workflows their teams use to get early user feedback on LLM-powered applications.
In the MobileMaker condition, users could use the NL revision feature to revise both the AI and UI functionality while on their phones. In the baseline condition, users could alter the prompt in a text editor. In both conditions, if users had trouble altering the prototype or if the NL revision did not work as expected (e.g., the revised prototype did not fit the user’s request), the study facilitators stepped in to help out.
The two prototypes participants experienced were both mobile, multimodal generative AI applications that two designers from our formative study were interested in exploring: (1) Img2Playlist, an application that generates a custom playlist based on an input image, and (2) Img2VideoIdeas, an application that generates ideas for videos to create, also based on an input image. We decided on these applications because they target a general audience of users, have the same core interaction pattern, and are actual mobile, generative AI applications professional designers would like to build.
@@ V-A Procedure
The overall outline for the exploratory user study is as follows: (1) Participants filled out a consent form prior to the study. (2) During the study, participants spent 50 minutes giving feedback on two different prototypes, one with MobileMaker (25 minutes) and the other with Figma and a prompt editor (25 minutes), all while thinking aloud. Condition-order and prototype-order were counterbalanced. (3) Participants then completed a post-study questionnaire that compared their experiences giving feedback on each prototype. (4) In a semi-structured interview, participants compared their experiences giving feedback in each condition. The total time commitment of the study was one hour.
To situate the task, in both conditions, users were asked to imagine they were the target user of both of these mobile applications; give any and all feedback they had while interacting with both prototypes; and alter the prototype when they had an applicable revision in mind. In both conditions, if they felt they hit a dead end with their revisions, users could restart from the beginning, either with the initial prototype (in the MobileMaker condition) or the initial prompt (in the baseline condition).
@@
With {Setup A/B}, I feel like I had a realistic or authentic experience of the envisioned application.
TABLE I: Post-task questionnaire filled out by participants after they gave feedback on the prototypes from both conditions. Each statement was rated on a 7-point Likert scale.
@@ V-C Questionnaire
The measures in the questionnaire (Table I) are based on prior literature that has established the qualities of a good prototype; we reformulated these qualities to fit the perspective of a user giving feedback. For instance, one quality of a prototype is its ability to capture the essence of the envisioned application. We thus measured if participants felt they had authentically experienced the envisioned application. Finally, to compare the ratings from the two conditions, we conducted paired sample Wilcoxon tests with full Bonferroni correction, since the study was within-subjects and the questionnaire data was ordinal.
@@ VI Findings
VI-A MobileMaker Supported Authentic Experimentation
Based on the questionnaire results, users felt that MobileMaker provided a significantly more authentic experience than the baseline prototype. This authenticity helped them (1) evaluate if the AI’s output was mobile appropriate, (2) experiment with serendipitous edge cases via unconventional, in-the-wild inputs, (3) discover discrepancies between their interpretation of the task and the model’s interpretation, and (4) more critically evaluate the model using contextual cues, leading to their identification of rich contextual input in their surroundings that the model did not have access to. One limitation of in situ testing was that testers couldn’t always easily get themselves into representative contexts (e.g., due to scheduling and location constraints), so it wasn’t always possible to quickly find the most “canonical” inputs (e.g., for Img2Playlist, a coffee shop or driving commute picture).
VI-A1 MobileMaker helped users evaluate if the product concept was mobile appropriate
While experiencing AI prototypes on MobileMaker, users discovered ways in which the prototypes may not be compatible with in-the-wild contexts or form factors. For example, P4 revised Img2VideoIdeas to produce a video script along with each video idea, but after trying a few different image inputs, they soon realized that “generating a full script for the video ideas on the go is overwhelming and too much text.” Beyond influencing considerations around form factor and mental capacity, the mobile vs. desktop context of use also affected user expectations around the overall “product fit” and appropriateness for mobile settings. For example, in the Img2VideoIdeas prototype, the model tended to output long-form content ideas, such as how-to videos and history-of videos. Though these may seem acceptable on desktop, P9 (who experienced Img2VideoIdeas in the MobileMaker condition) remarked that they expected more short-form, TikTok-style videos. In contrast, P16 (who experienced Img2VideoIdeas in the desktop condition) mentioned that the desktop setting may have prompted them to assume that the prototype was for creators making long-form content. Hence, by enabling participants to experience the prototype authentically on mobile, MobileMaker enabled users to not only assess if the look and feel of the model output was mobile-appropriate, but also to reflect on the overall product concept.
@@ VI-A2 Experiencing AI prototypes in the wild enabled serendipitous experimentation with edge cases
The photos that participants took in the wild tended to be “noisier” than on desktop, such as images taken at odd angles, images containing multiple objects of interest, images containing a combination of text and images, etc. The inherent noise in in-the-wild multimedia enabled users to probe edge cases and boundaries of the model’s capabilities and limitations. While testing Img2Playlist with MobileMaker, P1 noticed a picture of a dog on a nearby desk, with “Brandy” written at the edge. Curious to see if the model would pick up on this text, P1 input this photo, and to their surprise, the first song, “Brandy (You’re a Fine Girl)” by Looking Glass, specifically referenced the text. P1 noted that, with MobileMaker, “you are walking around, which allows you to think more about ideas and edge cases,” whereas on desktop, “I don’t think I would have gotten that [an image with text] just sitting here [using image search].” Similarly, mobile constraints also led to situations in which the specific content intended to be captured was not always obvious to the model. For instance, P6 took a picture of a far-away cherry blossom, where the photo did not center on the blossom and included other objects in the scene. They were surprised when the model produced a general playlist about New York rather than the cherry blossoms. Overall, encountering these edge-case scenarios helped users probe the boundaries of the model’s capabilities and limitations.
@@
In addition to encountering model edge cases (e.g., images with unique compositions), participants were also more likely to encounter objects and settings that were unique or unconventional given the use case (e.g., for Img2Playlist: pictures of hairspray, carton of milk, blank walls, etc.). This spontaneity and serendipity helped them uncover multiple possible interpretations of the problem. For example, while interacting with Img2Playlist on mobile, P2 input an image of a carton of milk, expecting a playlist with a “coffee shop vibe,” but instead received a more literal response: songs with titles related to milk (e.g., ”Cream” by Prince). These spontaneous experiences uncovered multiple interpretations of the problem: it is unclear if a “relevant” playlist means that it ought to be relevant to the literal object in the picture, to the vibe of the scene, or to the look of an artist’s album cover.
Meanwhile, on desktop, participants tended to search for conventional images that more often yielded expected model outputs. For example, P8 searched for “John Legend” in an image search engine and selected an image of one of his canonical album covers. They expected Img2Playlist to produce a playlist with songs only from that album, and it mostly did. This targeted approach lacked the serendipity of the MobileMaker inputs. As P8 explained, “If I do a search on Google, I have an intention already… there’s no randomness, and randomness is useful.” It’s possible that image search results may be less likely to expose intent mismatches between users and the AI, because they themselves were found through being highly linked to a user’s explicit text query. Overall, the images participants sourced in the baseline condition were more intent-driven and less ambiguous, which led to fewer discoveries of model output discrepancies.
@@ VI-A4 Contextual cues enabled a more informed evaluation of the model’s outputs
When experiencing prototypes on MobileMaker, participants had access to a wealth of sensory, spatial, and temporal context surrounding their image inputs, including the sounds and scents present in the environment, the encompassing landscape surrounding the image, and the events that occurred before and after the image was taken. This additional context helped them more critically evaluate model outputs that they would have otherwise deemed reasonable in the baseline condition. For example, P7 took a picture of two coworkers playing a relaxed game of pool in the early afternoon, when most people had just finished eating lunch. They were surprised when the model erroneously identified the scene as “competitive.” P7 realized that it may be difficult for a model to accurately assess the vibe with just a single, static image. “On the phone, I was in the environment in which the photo was taken… I have more context, not just what’s captured in that one-frame, but what’s also before or after. It makes me more critical about the result.” Similarly, P4 (while indoors) wished the model could take into consideration the fact that it was raining just outside their window, evoking a solemn mood. Overall, testing the prototype in-situ not only enabled users to assess the model’s outputs through the surrounding context it was situated in, but also helped them identify contextual cues the model did not have access to.
@@ VI-B How NL Revision Changed AI Prototyping Practices
Prompted by these in-the-wild discrepancies, participants revised the prototypes, ranging from adding UI controls (e.g. a dropdown menu for music genres or an input textbox to specify the audience for the video ideas) to changing model behavior (e.g. generating a description that associates each song in the playlist to the input image, or outputting a video script to accompany each idea). NL revision enabled them to play an active role in the design process, potentially shortening the test-feedback loop. However, user considerations around what types of revisions were actionable in MobileMaker may have constrained their ability to give higher-level feedback.
@@
VI-B1 Revising enabled users to provide better guardrails to the AI and reconsider the prototype’s interaction pattern
@@
When the model behavior did not match user intent, participants leveraged the NL revision tool to revise the prototype, such as by adding UI controls. As described in sec VI-A3, P2 discovered that the model sometimes generated playlists based on the literal object (e.g., “Cream” by Prince) rather than the vibe of the scene. When confronted with this issue, P8 revised their prototype to add a dropdown menu that allows users to select whether the model should generate playlists based on the “vibe” or the “object” shown in the image. In subsequent testing, they found that having this input helped guardrail the model from solely providing playlists that were literally related to the objects in the image. Similarly, when P6 took a far-away picture of cherry blossoms, and the model did not pick up on the cherry blossom in the cluttered image (described in sec VI-A2), P6 added a textbox input widget so that users can specify which object they wanted the playlist to be based on. After experimenting with this revision, P6 started to prefer having the playlist generated by the text. They appreciated the flexibility starting with text provided, and ultimately, they questioned the interaction pattern of starting with an image. Overall, the NL revision feature enabled participants to proactively react to model discrepancies, as well as critically reflect on the core interaction pattern (e.g., input and output types) of the application.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(a) Example desktop inputs
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(b) Example in-the-wild inputs
@@
Figure 5: Examples of test images tried by study participants during the desktop (5(a)) and in-the-wild (5(b)) conditions.
@@
VI-B2 NL revise enabled users to critically evaluate their own feedback in situ

### v2

@@ IV-C Prototype Testing Experience on Mobile
Designers can share their prototype via URL, which enables team members or other testers to try it out in the wild (DG2). When running a prototype on a mobile device, each set of user inputs, model output, and user feedback is saved as a test case. Designers can review all collected test cases in a testing dashboard, allowing them to quickly explore use cases and discover unexpected edge cases.
@@ IV-C1 Dynamic Prototype Revision
When testers experience unexpected outputs or prototype behavior, or they have feedback on how the prototype could be improved to better meet their needs, they can leverage the NL revision feature (sec. IV-B2) to generate a revised version of the prototype; this can be sent back to the prototype designer as a suggested revision. Each suggested revision includes the tester’s revision request, the updated JSON prototype specification, and the latest test case generated with the revised prototype prior to submission. After revising, testers can choose to either continue testing with the modified prototype or revert to the original specification, allowing for both continued iteration on a specific concept and exploration of multiple orthogonal ideas.
@@ IV-C2 Prototype Revision Dashboard
Prototype designers can use the revision dashboard (Fig. 4) to review and examine all proposed revisions, as well as dynamically try out each revision by applying its changes (Fig. 4A). When a revision is applied, MobileMaker renders the layout and functionality of its JSON prototype specification in the mobile phone preview, then populates the preview with inputs and outputs from the associated test case that the original tester generated and experienced. This enables prototype designers to quickly understand and explore the impact of each revision on their prototype, significantly enhancing and accelerating the prototype design loop (DG3).
Figure 4: Prototype Revision Dashboard
@@ V Exploratory User Study
Using MobileMaker as a probe, we conducted an exploratory study to understand: (1) how in the wild tester feedback on functional LLM-powered prototypes compares to status quo tester feedback collected on desktop and (2) how revising in the wild affects tester feedback.
The study was within-subjects with two conditions. In each condition, participants (henceforth referred to as “testers”) were asked to give feedback on a prototype. In the MobileMaker condition, testers gave feedback on a MobileMaker prototype in the wild, and they were also able to revise the prototype with the NL revision feature. In the baseline (desktop) condition, testers clicked through a Figma prototype (with the same UI as the MobileMaker prototype) that demonstrated the basic interaction. Afterwards, they searched for images online to test the LLM prompt that would power the prototype in the Google AI Studio prompt editor. This baseline setup was informed by our formative studies with designers, who indicated that these are typical prototyping workflows their teams use to get early tester feedback on LLM-powered applications.
In the MobileMaker condition, testers used the NL revision feature to revise both the AI and UI functionality while on their phones. In the baseline condition, testers could alter the prompt in a text editor. In both conditions, if testers had trouble altering the prototype or if the NL revision did not work as expected (e.g., the revised prototype did not fit the tester’s request), the study facilitators stepped in to help out.
The two prototypes testers experienced were both mobile, multimodal generative AI applications that two designers (D1&2) from our formative study ideated: (1) Img2Playlist, an application that generates a custom playlist based on an input image, and (2) Img2VideoIdeas, an application that generates ideas for videos to create, also based on an input image. Each tester experienced both prototypes, one per condition and counterbalanced (e.g. P1 experienced Img2Playlist with MobileMaker and Img2VideoIdeas with the baseline, whereas P4 experienced Img2Playlist with the baseline and Img2VideoIdeas with MobileMaker). We decided on these applications because they target a general audience of users, have the same core interaction pattern, and are actual mobile, generative AI applications professional designers would like to build.
@@ V-A Procedure
The overall outline for the study is as follows: (1) Testers filled out a consent form prior to the study. (2) During the study, testers spent 50 minutes giving feedback on two different prototypes, one with MobileMaker (25 minutes) and the other with Figma and a prompt editor (25 minutes), all while thinking aloud. Condition-order and prototype-order were counterbalanced. (3) Testers then completed a post-study questionnaire that compared their experiences giving feedback on each prototype. (4) In a semi-structured interview, testers compared their experiences giving feedback in each condition. Each study session took one hour on average. To situate the task, in both conditions, testers were asked to imagine they were the target user of both of these mobile applications; give any and all feedback they had while interacting with both prototypes; and alter the prototype when they had an applicable revision in mind. In both conditions, if they felt they hit a dead end with their revisions, testers could restart from the beginning, either with the initial prototype (in the MobileMaker condition) or the initial prompt (in the baseline condition).
@@
With {Setup A/B}, I felt like I had a realistic or authentic experience of the envisioned application.
TABLE I: Post-task questionnaire filled out by testers after they gave feedback on the prototypes from both conditions.
@@ V-C Questionnaire
The questionnaire measures (Table I) are derived from established literature on qualities of good [66, 67, 68] and have been adapted to fit the perspective of testers. For instance, one quality of a prototype is its ability to capture the essence of the envisioned application [67]. We thus measured if testers felt they had authentically experienced the envisioned application. Finally, to compare the ratings from the two conditions, we conducted paired sample Wilcoxon tests with full Bonferroni correction, since the study was within-subjects and the questionnaire data was ordinal.
@@ VI Findings
VI-A How Testing in the Wild Impacted Tester Feedback
Based on the questionnaire results, testers felt that MobileMaker provided a significantly more authentic experience than the baseline prototype, because they could interact with a functional prototype in the wild. Experiencing a functional prototype in the wild helped testers: (1) evaluate if the AI’s output was mobile appropriate, (2) experiment with serendipitous edge cases via unconventional, in-the-wild inputs, (3) discover discrepancies between their interpretation of the task and the model’s interpretation, and (4) more critically evaluate the model using contextual cues, leading to their identification of rich contextual input in their surroundings that the model did not have access to. One limitation of in situ testing was that testers couldn’t always easily get themselves into representative contexts (e.g., due to scheduling and location constraints), so it wasn’t always possible to quickly find the most “canonical” inputs (e.g., for Img2Playlist, a coffee shop or driving commute picture).
VI-A1 MobileMaker helped testers evaluate if the product concept was mobile appropriate
While experiencing AI prototypes on MobileMaker, testers discovered ways in which the prototypes may not be compatible with in-the-wild contexts or form factors. For example, P4 revised Img2VideoIdeas to produce a video script along with each video idea, but after trying a few different image inputs, they soon realized that “generating a full script for the video ideas on the go is overwhelming and too much text.” Beyond influencing considerations around form factor and mental capacity, the mobile vs. desktop context of use also affected user expectations around the overall “product fit” and appropriateness for mobile settings. For example, in the Img2VideoIdeas prototype, the model tended to output long-form content ideas, such as how-to videos and history-of videos. Though these may seem acceptable on desktop, P9 (who experienced Img2VideoIdeas in the MobileMaker condition) remarked that they expected more short-form, TikTok-style videos. In contrast, P16 (who experienced Img2VideoIdeas in the desktop condition) mentioned that the desktop setting may have prompted them to assume that the prototype was for creators making long-form content. Hence, by enabling testers to experience the prototype authentically on mobile, MobileMaker enabled testers to not only assess if the look and feel of the model output was mobile-appropriate, but also to reflect on the overall product concept.
@@ VI-A2 Experiencing AI prototypes in the wild enabled serendipitous experimentation with edge cases
The photos that testers took in the wild tended to be “noisier” than on desktop, such as images taken at odd angles, images containing multiple objects of interest, images containing a combination of text and images, etc. The inherent noise in in-the-wild multimedia enabled testers to probe edge cases and boundaries of the model’s capabilities and limitations. While testing Img2Playlist with MobileMaker, P1 noticed a picture of a dog on a nearby desk, with “Brandy” written at the edge. Curious to see if the model would pick up on this text, P1 input this photo, and to their surprise, the first song, “Brandy (You’re a Fine Girl)” by Looking Glass, specifically referenced the text. P1 noted that, with MobileMaker, “you are walking around, which allows you to think more about ideas and edge cases,” whereas on desktop, “I don’t think I would have gotten that [an image with text] just sitting here [using image search].” Similarly, mobile constraints also led to situations in which the specific content intended to be captured was not always obvious to the model. For instance, P6 took a picture of a far-away cherry blossom, where the photo did not center on the blossom and included other objects in the scene. They were surprised when the model produced a general playlist about New York rather than the cherry blossoms. Overall, encountering these edge-case scenarios helped testers probe the boundaries of the model’s capabilities and limitations.
@@
In addition to encountering model edge cases (e.g., images with unique compositions), testers were also more likely to encounter objects and settings that were unique or unconventional given the use case (e.g., for Img2Playlist: pictures of hairspray, carton of milk, blank walls, etc.). This spontaneity and serendipity helped them uncover multiple possible interpretations of the problem. For example, while interacting with Img2Playlist on mobile, P2 input an image of a carton of milk (Figure 6b), expecting a playlist with a “coffee shop vibe,” but instead received a more literal response: songs with titles related to milk (e.g., ”Cream” by Prince). These spontaneous experiences uncovered multiple interpretations of the problem: it is unclear if a “relevant” playlist means that it ought to be relevant to the literal object in the picture, to the vibe of the scene, or to the look of an artist’s album cover.
Meanwhile, on desktop, testers tended to search for conventional images that more often yielded expected model outputs. For example, P8 searched for “John Legend” in an image search engine and selected an image of one of his canonical album covers. They expected Img2Playlist to produce a playlist with songs only from that album, and it mostly did. This targeted approach lacked the serendipity of the MobileMaker inputs. As P8 explained, “If I do a search on Google, I have an intention already… there’s no randomness, and randomness is useful.” It’s possible that image search results may be less likely to expose intent mismatches between testers and the AI, because they themselves were found through being highly linked to a user’s explicit text query. Overall, the images testers sourced in the baseline condition were more intent-driven and less ambiguous, which led to fewer discoveries of model output discrepancies.
@@ VI-A4 Contextual cues enabled a more informed evaluation of the model’s outputs
When experiencing prototypes on MobileMaker, testers had access to a wealth of sensory, spatial, and temporal context surrounding their image inputs, including the sounds and scents present in the environment, the encompassing landscape surrounding the image, and the events that occurred before and after the image was taken. This additional context helped them more critically evaluate model outputs that they would have otherwise deemed reasonable in the baseline condition. For example, P7 took a picture of two coworkers playing a relaxed game of pool in the early afternoon (Figure 6b), when most people had just finished eating lunch. They were surprised when the model erroneously identified the scene as “competitive.” P7 realized that it may be difficult for a model to accurately assess the vibe with just a single, static image. “On the phone, I was in the environment in which the photo was taken… I have more context, not just what’s captured in that one-frame, but what’s also before or after. It makes me more critical about the result.” Similarly, P4 (while indoors) wished the model could take into consideration the fact that it was raining just outside their window, evoking a solemn mood. Overall, testing the prototype in-situ not only enabled testers to assess the model’s outputs through the surrounding context it was situated in, but also helped them identify contextual cues the model did not have access to.
@@ VI-B How NL Revision Changed AI Prototyping Practices
Prompted by these in-the-wild discrepancies, testers revised the prototypes, ranging from adding UI controls (e.g. a dropdown menu for music genres or an input textbox to specify the audience for the video ideas) to changing model behavior (e.g. generating a description that associates each song in the playlist to the input image, or outputting a video script to accompany each idea). NL revision enabled testers to play an active, fulfilling role in the design process, and it also shortened the test-feedback loop. However, revising also potentially limited tester feedback to smaller, incremental revisions of the application’s functionality. Whereas with Figma, testers gave higher level feature requests and ideas. This suggests that perhaps the feedback these prototypes provide are complementary and useful at different stages of the design process.
@@
VI-B1 Revising enabled testers to provide better guardrails to the AI and reconsider the prototype’s interaction pattern
@@
When the model behavior did not match user intent, testers leveraged the NL revision tool to revise the prototype, such as by adding UI controls. As described in sec VI-A3, P2 discovered that the model sometimes generated playlists based on the literal object (e.g., “Cream” by Prince) rather than the vibe of the scene. When confronted with this issue, P8 revised their prototype to add a dropdown menu that allows testers to select whether the model should generate playlists based on the “vibe” or the “object” shown in the image. In subsequent testing, they found that having this input helped guardrail the model from solely providing playlists that were literally related to the objects in the image. Similarly, when P6 took a far-away picture of cherry blossoms, and the model did not pick up on the cherry blossom in the cluttered image (described in sec VI-A2), P6 added a textbox input widget so that testers can specify which object they wanted the playlist to be based on. After experimenting with this revision, P6 started to prefer having the playlist generated by the text. They appreciated the flexibility starting with text provided, and ultimately, they questioned the interaction pattern of starting with an image. Overall, the NL revision feature enabled testers to proactively react to model discrepancies, as well as critically reflect on the core interaction pattern (e.g., input and output types) of the application.
Example desktop inputs
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Example in-the-wild inputs
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 6: Examples of test images tried by study participants during the desktop (a) and in-the-wild (b) conditions.
@@
VI-B2 NL revise enabled testers to critically evaluate their own feedback in situ